UT Nov 20 2021 - D/E Night shared with PEPSI/PETS



Night Info

  • Sunset 17:22
  • Twilight Ends 18:42
  • Handover to PEPSI 22:45
  • Handover from PEPSI to LBTO 03:01
  • Twilight Begins 05:29
  • Sunrise 06:49

Night Plan

  • Starting config bino-LUCI
  • LUCI-1 AO/ESM Testing
    • Full AO:
      • OSCO
      • thin clouds (up to 1mag extinction), seeing 1.2" or better (1.5" or better for ESM)
      • OT file: /home/lbto/Service_Observing/2021B_D_E/LUCI/AO_ESM/AO/AO_Testing_V31.xml
      • Close the loop on BIn1
      • Close the loop on Bin2
    • Close the loop on Bin 4/ESM targets
      • ESM Script Location: /home/lbto/Service_Observing/2021B_D_E/LUCI/AO_ESM/ESM/ESM_V25.xml
        • Scripts in M52 sub-directory
        • M52 R=15.2 (script forces R=16.51 to make it Bin4)
        • Option for LUCI-1 only or mixed-mode
        • Cycles through K,H,J - 10min per filter exposure time - dithers out of field
        • Scripts already generated for mixed-mode with corrections
  • M31 Nova - if time after AO -OR- if AO halts for technical reasons
    • File Location: /home/lbto/Service_Observing/2021B_D_E/MODS/RMWagner_M31_Nova
    • Acq Image time 120-150sec
    • Exposures 4x1800sec Dual Grating
  • Reconfigure for PEPSI 22:45hr
  • If more AO testing needed - reconfigure from PEPSI to bino-LUCI 03-03:15hr
  • Otherwise - reconfigure for bino-MODS
  • GRB211023A followup - Dual Grating - 2hrs w/overheads
    • thin clouds (up to 0.5mag extinction), seeing 1.2" or better on the DIMM
    • File Location: /home/lbto/Service_Observing/2021B_D_E/MODS/Science_Rothberg/GRB211023A
    • File: ot_rossi_grb211023A-sn_mods_V2.xml
    • Finder image from first MODS-1R acquisition images (example for on-target and blind offset view - orientation matches script)
    • Option for target acquisition or blind offset (star would be shifted to the segment just above the target)
    • Exposures are 3x1800sec
  • SpecStd G191B2B - take during twilight if time

  • Poor Conditions Backup:
    • Bino-LUCI
    • Local Mergers
      • OT File: /home/lbto/Service_Observing/2021B_D_E/PoorWeather/LocalMergers/COBandhead_Mergers_V29.xml
      • Any conditions (no priority order)
      • NGC 6598 - 1hr bino (repeat script 1x) if really bad, go longer - twilight thru 20:30hr
      • DONE NGC 7585 - 1hr bino (repeat script 1x) if really bad, go longer - twilight until 23hr
      • DONE UGC 11905 - 1hr bino (repeat script 1x) if really bad, go longer - twilight until 23hr
      • Arp 230 - 1hr bino (repeat script 1x) if really bad, go longer - 18:30hr thru handover
        • telluric rises 19:30hr thru handover
        • DONE One single pair on Arp 230 observed plus telluric before PEPSI handoff.
      • NGC 4004 - 1hr bino (repeat script 1x) if really bad, go longer - 03:30 - twilight

Summary

LUCI AO setup was attempted at the start of the night, but there was a critical failure with the homing of the pupil rotator. See IT8524 for details. Thick clouds moved in so poor weather LUCI programs were opted for once LUCI AO was passed on. Conditions improved around 4:15 UT, when the clouds appears to dissipate before hitting the summit for the most part and seeing stabilized. At this point it was too late to switch to mods before handoff to PEPSI at 5:45UT. PEPSI PETS was executed, starting target observations from 6UT through 10UT. Reconfigured to MODS for GRB211023. Seeing brecame highly variable during the GRB observation with periods up to 2", but mostly around 1.1-1.2"

Overview

People of the Night: Jenny, Doug, Barry, Josh, Ilya

23:40 LUCI started and several 2.51 test images taken. Tried to start up MODS on robs but got the following persistently when bringing up modsDisp:

Launching MODS1R ds9 display, please wait...
Traceback (most recent call last):
File "/lbt/lbto/mods/bin/modsDisp", line 375, in <module>
md = start_ds9(ds9ID, timeout=30)
File "/lbt/lbto/mods/bin/modsDisp", line 228, in start_ds9
raise RuntimeError (f"failed to detect any ds9 window "
RuntimeError: failed to detect any ds9 window with title 'MODS1R-1637365516-3048735'

Previous image shows none in mods1 dashboard, likely from lastnights ssh drop. Taking a test image before brining up modsDisp. mods Disp was fine after the simSnap.

Josh executed wfs_init but that failed
He called Juan Carlos and executed a w_stop_all, w_start_all
Josh had a failure bringing up the WFS. Timeout on the pupil rotator. Juan Carlos suggested a retry which failed. Then he suggested the homing the pupil rotator which is a long endevor. Josh is executing.

0:16 The pupil homing has been going on for 10 + minutes and is moving a degree every 10 sec which is abnomally slow. Josh tried to abort after Doug directed him to, but no luck. Its still going. Josh is preparing to open.

0:19 Doug is trying to power cycle the pupil rotator and giving it a kick. But that failed. The pupil rotator is continuing to try and find home at 1 degree/10sec.

0:32. The pupil rotator appears to have gone around more than once without finding home. Doug is fully powering off the wfs to see if the pupil rotator issues are caused by a communication issue. w_stop_all, w_start_all

0:43 The WFS power cycle does not seem to have resolved the issue. The homing is still failing. Juan Carlos (who is not on tonight but did test during the day) just submitted an IT on this #8524 so surprise!

1:14 The pupil rotator has made 2 full rotations without finding home. We are going to authorize bino LUCi and execute Poor weather LUCI programs. We have a 1-2 mags extinction from clouds.

1:20 Josh is having a problem with authorizing luci2.

1:30 LUCI2 MOS error, Storage to FPU failed for 0.5" longslit (#7), MHU translator translator failed to reach 0.5" longslit mask position withing tolerance
SG off, grabber neg lim, Rotator 340.988
Spindle 274.966, -4757 ** Actual 275.5+/-4.0, -4748+/-5

Checking preconditions for Init OK, Init Ok back in buisiness

1:37 Josh checking pointing and collimation

1:44 Sending preset for telluric for NGC 7585 - HD21807. Seeing is 0.67" on the guiders
Luci1 offset -10.4818, -0.2156
Luci2 offset -10.3078, 0.0226

1:54 Science started, 0.64" on teh guiders, spotted clouds overhead

luci1.20211120.0024-26
luci2.20211120.0010-12

The folks are all leaving us, just josh and me, and Juan carlos has joined.

1:59 Preset to NGC7585.

Luci1 offset -0.7351, -0.6232, additional 0.14 needed for luci1
Luci2 offset -0.3142, -0.9984
When doing the Longslit alignment the Alignment and Background images did not auto populate.

2:16 starting science for NGC7585. seeing 0.7" on teh guiders and cirrus blanketting the sky.

luci1.20211120.033
luci2.29211129.017

2:35 Seeing 0.65" on the guiders, in a "clearing" with negligeable extinction, but a thick band of cloud is moving in.

2:39 Thick clouds are here, lost guiding. Pausing script so we don't offset while we don't have a guide star.
clouds moved in at the tail end of:

luci1.20211120.036
luci2.29211129.020

2:44 Somehow our guide star has returned so we are resuming, although there are some pretty opaque looking clouds overhead. This may be short lived, hovering about 1-1.5 mags of extinction.

luci1.20211120.037
luci2.29211129.021

2:49 Doug is homing the AO stages so there may be some movement of the stages in the lower 20-30 columns in luci1.20211120.037. Cube rotator is still spinning, doug will continue to try int eh background and submit and update to the IT.

3:08 A few dips periodically of 1+ mags of extinction but for the most part we are in a "clearing" with relatively thin cirrus and <0.5 mag extinction. Seeing has jumped a bit up to 1" on the guiders, and is starting to vary up to 1.3 periodically. More thick bands looming on the horizon.

3:18 Seeing bounding between 0.7-1.3" on the guiders. Thick clouds heading in, although just cirrus overhead for the moment. Tonight does not look suitable to switch to MODS for the Nova, especially since we would have to do so about now. Given the clouds, and now variable seeing we will remain with the poor weather backups.

3:22 Seeing was good for the majority of exposures for NGC 7585, averaging 0.7", and for the most part <1mag extinction (paused when it was really bad). I will keep it at only one run through the script and move on to UGC 11905

luci1.20211120.033-44
luci2.29211129.017-28

3:31 Doug has finished his attempts with the WFS unit and has fully powered it off. He has submitted an update to the IT. Slewing to telluric for UGC 11905 - BD +14. Left failed to find a guide star so Josh is executing pointing check

3:37 Resending preset to telluric for UGC11905

luci1 offset -10.7893, 0.8988
luci2 offset -10.4209, -0.1559

3:45 starting telluric science seeing bouncing all over 0.6-1.2". thin cirrus overhead with thick clouds looming on the horizon and on the satellite.

luci1.20211120.048-50
luci2.29211129.032-34

3:51 Sending preset to UGC11905

luci1 offset -0.7021, -0.5878
luci2 offset 0.7282, -0.4474, -0.05 in X needed

4:06 Seeing 0.65-0.75" on the guider, thick blobby (yes that is the technical term) clouds are moving in.

luci1.20211120.055-66
luci2.29211129.040-51

4:15 The thick cloud was short lived and we are in a clearing with thin cirrus. We had 4 mags of extinction briefly. Thick clouds looming on the horizon. Seeing appears to have stabilized again to 0.65" on the guiders.

4:53 The satellite shows the clouds breaking as they approach the summit. We have remained with light cirrus and seeing 0.65-0.8" since ~4:15 although there are heavier clouds along the horizon

5:07 some thicker clouds are starting to move in. Seeing ~ 1mag of extinction. Only 2 exposures remaining for this target.

5:16 Sending preset to Arp230 Telluric. We dont have enough time to get much on this target but can get a few frames.
luci1 offset -10.4677, -1.3747
luci2 offset -10.5196, -0.7759

5:23 Started science on Arp 230 telluric, thin cirrus overhead. Seeing 0.65-1.0" on guider.

luci1.20211120.070-72
luci2.29211129.055-57

5:31 Taking one pair set on Arp230, cirrus, seeing 0.65-0.8"
luci1 offset -0.3204, -0.5584
luci2 offse1 -0.6724, -2.9131

luci1.20211120.0077-78
luci2.20211120.0062-63

5:47 LUCI safed and reauthorizing for PEPSI.

5:52 Pointing and collimation preset near pepsi PETS target.

6:02 Ilya has started exposuing on PETS target. Thin cirrus and good seeing throughout.

10:03 Gave Ilya 3 extra minutes to finish his last exposure. Josh was ready with his finger on the button ready to reconfig to MODS.

10:17 Sending pointing and collimation near GRB

10:26 Sendig preset to GRB211023A. Used blind offset script since we do have bright sky and in the interest of efficiency. Thin cirrus overhead, seeing ~1-1.1" on the guiders, 0.8" on teh DIMM

mods1 offset -1.379, 11.436, additional -0.068, 0.144, then blind offset

mods2 offset 3.062, 9.349, then blind offset

Source faintly vsible in mods2 60sec so likely would have been ok with source acquisition with 120sec, but blind offset scertainly more efficient. Note that the science scripts were set upt with 3x1800 at 2 dither positions (total 6x1800). We wont be able to complete that, but above suggests that may bhave been a script error, reminant of a past execution.

10:42 Starting science

IMCS lock failed on both sides, but retry worked. Seeiing is showing 1.4" on DX periodically, but 1.1" for the most part. SX 1-1.1". DIMM 0.9-1"

mods1r.20211120.0007-9, mods1b.20211120.0003-5
mods2r.20211120.0006-8, mods2b.20211120.0003-5

11:30 The traces is very very faint. Seeing has been around 1.2" on the left and 1.3" on the right

11:50 Seeing is deteriorating, 1.1-1.4" on the left, 1.2-1.5" on the right, mostly clear. 3/3 exposure started

12:09. Seeing has completely deteriorated. 1.25-2", highly variable. mostly clear overhead.

12:20 Starting a 4th exposure to collect a few more photons. Seeing seems to be bouncing around but is coming back down again, 1-1.2" on the left, 1.2-1.35" on the right

12:25 I spoke too soon, still variable, and bouncing around with periods up to 2", average closer to 1.2-1.3"

12:38 Stopping and reading out the shortened obs
mods1r.20211120.0010, mods1b.20211120.0006
mods2r.20211120.0009, mods2b.20211120.0006

12:40 heading to G191B2B. G191B2B is very close to the moon.... Going to GD71 instead. Not ideal, I had planned for the shorter times with G191B2B. Whoops.

mods1 keeps failing with inst config already in progress. Update resolved.

12:51 Starting mods2 science after imcs lock fail while mods1 catches up

12:56 mods1 required and additional offset , imcs lock fail, to boot so finally started.

mods1r.20211120.0013-14, mods1b.20211120.0007-8
mods2r.20211120.0012-14, mods2b.20211120.0007-9

13:10 sky is high in these exposures, so may not be all that useful. Stopping at only 2 for mods1 and letting Josh close up.

Calibrations taken:
N1.8 G210 K@2.27 LS 0.5
Counts seem low for flats... Will redo. Flat times were not scaled for 0.5" slit width. These are older scripts so perhaps there is a reason. Tried a few combinations, running with the halo1 with luci1 was more efficient.

N1.8 G210 K@2.3 LS 0.5 - taken with original flat times and adjusted

mods dual grating 5" flats

-- BarryRothberg - 19 Nov 2021
Topic revision: r5 - 20 Nov 2021, JenniferPower
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback